OpenStudy (fernandez555):

determine the equation of a line in slope intercept form for points b (5,5) and c (1,3) -show steps in addition to the final equation please

OpenStudy (jidnyesha):

m=(y2-y1)/(x2-x1) m=(3-5)/(1-5) m=(-2)/(-4) m=.5 y=mx+c so taking (5,5) y=5 and x=5 5=(.5*5)+c 5=2.5+c c=5-2.5 c=2.5 therefore, y=.5x+2.5

OpenStudy (anonymous):

slope is wrong

OpenStudy (pratyush5):